# On the Statistical Complexity of Estimating VENDI Scores from Empirical Data

### v1

@@ 6 Numerical Results
We evaluated the convergence of the VENDI score, the truncated VENDI score, and the proxy VENDI scores using the Nyström method and FKEA in our numerical experiments. We provide a comparative analysis of these scores across different data types and models, including image, text, and video. In our experiments, we considered the cosine similarity kernel as a standard kernel function with a finite-dimension map and the Gaussian (RBF) kernel as a kernel function with an infinite-dimension feature map. In the experiments with Gaussian kernels, we matched the kernel bandwidth parameter with those chosen by [2, 4] for the same datasets. We used 20,000 number of samples per score computation, consistent with standard practice in the literature. To investigate how computation-cutting methods compare to each other, in the experiments we matched the truncation parameter $t$ of our defined $t$-truncated VENDI score with the Nyström method’s hyperparameter on the number of randomly selected rows of kernel matrix and the FKEA’s hyperparameter of the number of random Fourier features. The VENDI and FKEA implementations were adopted from the corresponding references’ GitHub webpages, while the Nyström method was adopted from the scikit-learn Python package.
@@ 6.1 Convergence Analysis of VENDI scores
To assess the convergence of the discussed VENDI scores, we conducted experiments on four datasets including ImageNet and FFHQ[24] image datasets, a synthetic text dataset with 400k paragraphs generated by GPT-4 about 100 randomly selected countries, and the Kinetics video dataset [25]. Our results, presented in Figure LABEL:VENDI_convergence, show that for the finite-dimension cosine similarity kernel the VENDI score converges rapidly to the underlying value and the proxy versions including truncated and Nyström VENDI scores were almost identical to the original VENDI score. This observation is consistent with our theoretical results on the convergence of VENDI scores under finite-dimension kernel maps. On the other hand, in the case of infinite dimension Gaussian kernel, we observed that the $\mathrm{VENDI}_{1}$ score did not converge using 20k samples and the score value kept growing with a considerable rate. However, the $t$-truncated VENDI score with $t=10000$ converged to its underlying statistic shortly after 10000 samples were used. Consistent with our theoretical result, the proxy Nyström and FKEA estimated scores with their rank hyperparameter matched with $t$ also converged to the limit of the truncated VENDI scores. The numerical results show the connection between the truncated VENDI statistic and the existing kernel-based algorithms for approximating the VENDI score.
6.2 Correlation between the truncated VENDI statistic and diversity of data
@@ 7 Conclusion
In this work, we investigated the statistical convergence properties of VENDI diversity scores estimated from empirical samples. We highlighted that, due to the high computational complexity of the score for datasets larger than a few tens of thousands of generated data points, the score is often calculated using sample sizes below 20,000. We demonstrated that such limited sample sizes do not pose a problem for statistical convergence as long as the kernel feature dimension is bounded. Conversely, our numerical results showed a lack of convergence to the VENDI statistic when using an infinite-dimensional kernel map, such as the Gaussian kernel. To address this issue, we introduced the truncated VENDI statistic as an alternative target statistic. We further showed that existing methods, such as Nyström and FKEA, for approximating VENDI scores concentrate well around this truncated VENDI statistic. A promising direction for future research is to explore the relationship between other kernel approximation techniques and the truncated VENDI statistic. Moreover, a comprehensive analysis of the computational-statistical trade-offs involved in estimating the VENDI score remains an open question for future work.
@@ A.1 Proof of Theorem 1
To prove the theorem, we will use the following lemma followed from [28, 29].
Lemma 1 (Vector Bernstein Inequality [28, 29]).
@@
On the other hand, note that $\phi(x)\otimes\phi(x)$ is the vectorized version of rank-1 matrix $\phi(x)\phi(x)^{\top}$, which shows that the above inequality is equivalent to
$\mathbb{P}\Bigl{(}\Bigl{\|}\frac{1}{n}\sum_{i=1}^{n}\bigl{[}\phi(x_{i})\phi(x_{i})^{\top}\bigr{]}-\mathbb{E}_{x\sim P}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}\Bigr{\|}_{F}\geq\epsilon\Bigr{)}\,\leq\,\exp\Bigl{(}\frac{8-n\epsilon^{2}}{32}\Bigr{)}$
@@
Subsequently, we can apply the Hoffman-Wielandt inequality which shows that for the sorted eigenvalue vectors of $C_{X}$ (denoted by $\widehat{\boldsymbol{\lambda}}_{n}$ in the theorem) and $\widetilde{C}_{X}$ (denoted by $\widetilde{\boldsymbol{\lambda}}$ in the theorem) we will have $\|\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\|_{2}\leq\|C_{X}-\widetilde{C}_{X}\|_{F}$, which together with the previous inequality leads to
@@ A.2 Proof of Corollary 1
A.3 Proof of Corollary 2
Given $21\leq\alpha<$ 2, the inequality $\|v\|_{\alpha}\leq d^{\frac{2-\alpha}{2}}\|v\|_{2}$ holds for every $d$-dimensional vector $v\in\mathbb{R}^{d}$. Therefore, we can repeat the proof of Corollary 1 to show the following for every $21\leq\alpha<$ 2
@@ A.4 Proof of Theorem 2
Proof.
To prove the lemma, first note that $\mathbf{v}^{(t)}\in\Delta_{t}$, i.e. its first $t$ entries are non-negative and add up to $1$, and also its last $d-t$ entries are zero. Then, consider the projection problem discussed in the lemma:
@@
Then, since we know from the assumptions that $v_{i}\geq 0$ and $\sum_{i=1}^{t}v_{i}\leq 1$, the discussed $\mathbf{u}^{*}\in\mathbb{R}^{t}$ where $u^{*}_{i}=v_{i}+(1-S_{t})/t$ together with Lagrangian coefficients $\mu_{i}=0$ (for inequality constraint $u_{i}\geq 0$) and $\lambda=(1-S_{t})/t$ (for equality constraint) satisfy the KKT conditions. The primal and dual feasibility conditions as well as the complementary slackness clearly hold for these selection of primal and dual variables. Also, the KKT stationarity condition is satisfied as for every $i$ we have $u^{*}_{i}-v_{i}-\lambda-\mu_{i}=0$. Since the optimization problem is a convex optimization task with affine constraints, the KKT conditions are sufficient for optimaility which proves the lemma. ∎
Based on the above lemma, the eigenvalues $\widehat{\boldsymbol{\lambda}}_{n}^{(t)}$ used to calculate the $t$-truncated VENDI score $\mathrm{VENDI}^{(t)}_{\alpha}(x_{1},\ldots,x_{n})$ are the projections of the top-$t$ eigenvalues in $\widehat{\boldsymbol{\lambda}}_{n}$ for the original score $\mathrm{VENDI}_{\alpha}(x_{1},\ldots,x_{n})$ onto the $t$-simplex subset of $\mathbb{R}^{d}$ according to the $\ell_{2}$-norm. Similarly, the eigenvalues $\widetilde{\boldsymbol{\lambda}}_{n}^{(t)}$ used to calculate the $t$-truncated VENDI statistic $\mathrm{VENDI}^{(t)}_{\alpha}(P_{X})$ are the projections of the top-$t$ eigenvalues in $\widetilde{\boldsymbol{\lambda}}$ for the original VENDI statistic $\mathrm{VENDI}_{\alpha}(P_{x})$ onto the $t$-simplex subset of $\mathbb{R}^{d}$.

### v3

@@ 6 Numerical Results
Figure 3: Statistical convergence of Vendi score for different sample sizes on FFHQ[Karras et al., 2019] data: (Left plot) finite-dimension cosine similarity kernel (Right plot) infinite dimension Gaussian kernel with bandwidth $\sigma=35$. DINOv2 embedding (dimension 768) is used in computing the scores.
Figure 4: Statistical convergence of Vendi score for different sample sizes on Synthetic Countries data: (Left plot) finite-dimension cosine similarity kernel (Right plot) infinite dimension Gaussian kernel with bandwidth $\sigma=0.6$. text-embedding-3-large embedding (dimension 3072) is used in computing the scores.
Figure 5: Statistical convergence of Vendi score for different sample sizes on Kinetics400[Kay et al., 2017] data: (Left plot) finite-dimension cosine similarity kernel (Right plot) infinite dimension Gaussian kernel with bandwidth $\sigma=4.0$. I3D embedding (dimension 1024) is used in computing the scores.
Figure 6: Diversity evaluation of Vendi scores on truncated StyleGAN3 generated FFHQ dataset with varying truncation coefficient $\psi$. Fixed sample size $n=$20k is used for estimating the scores.
Figure 7: Diversity evaluation of Vendi scores on truncated StyleGAN-XL generated ImageNet dataset with varying truncation coefficient $\psi$. Fixed sample size $n=$20k is used for estimating the scores.
We evaluated the convergence of the Vendi score, the truncated Vendi score, and the proxy Vendi scores using the Nyström method and FKEA in our numerical experiments. We provide a comparative analysis of these scores across different data types and models, including image, text, and video. In our experiments, we considered the cosine similarity kernel as a standard kernel function with a finite-dimension map and the Gaussian (RBF) kernel as a kernel function with an infinite-dimension feature map. In the experiments with Gaussian kernels, we matched the kernel bandwidth parameter with those chosen by [Jalali et al., 2023, Ospanov et al., 2024b] for the same datasets. We used 20,000 number of samples per score computation, consistent with standard practice in the literature. To investigate how computation-cutting methods compare to each other, in the experiments we matched the truncation parameter $t$ of our defined $t$-truncated Vendi score with the Nyström method’s hyperparameter on the number of randomly selected rows of kernel matrix and the FKEA’s hyperparameter of the number of random Fourier features. The Vendi and FKEA implementations were adopted from the corresponding references’ GitHub webpages, while the Nyström method was adopted from the scikit-learn Python package.
@@ 6.1 Convergence Analysis of Vendi Scores
To assess the convergence of the discussed Vendi scores, we conducted experiments on four datasets including ImageNet and FFHQ [Karras et al., 2019] image datasets, a synthetic text dataset with 400k paragraphs generated by GPT-4 about 100 randomly selected countries, and the Kinetics video dataset [Kay et al., 2017]. Our results, presented in Figures 3, 4, and 5, show that for the finite-dimension cosine similarity kernel the Vendi score converges rapidly to the underlying value and the proxy versions including truncated and Nyström Vendi scores were almost identical to the original Vendi score. This observation is consistent with our theoretical results on the convergence of Vendi scores under finite-dimension kernel maps. On the other hand, in the case of infinite dimension Gaussian kernel, we observed that the $\mathrm{Vendi}_{1}$ score did not converge using 20k samples and the score value kept growing with a considerable rate. However, the $t$-truncated Vendi score with $t=10000$ converged to its underlying statistic shortly after 10000 samples were used. Consistent with our theoretical result, the proxy Nyström and FKEA estimated scores with their rank hyperparameter matched with $t$ also converged to the limit of the truncated Vendi scores. The numerical results show the connection between the truncated Vendi score and the existing kernel methods for approximating the Vendi score.
6.2 Correlation between the truncated Vendi score and diversity of data
@@ 7 Conclusion
In this work, we investigated the statistical convergence behavior of Vendi diversity scores estimated from empirical samples. We highlighted that, due to the high computational complexity of the score for datasets larger than a few tens of thousands of generated data points, the score is often calculated using sample sizes below 10,000. We demonstrated that such restricted sample sizes do not pose a problem for statistical convergence as long as the kernel feature dimension is bounded. However, our numerical results showed a lack of convergence to the population Vendi when using an infinite-dimensional kernel map, such as the Gaussian kernel. To address this gap, we introduced the truncated population Vendi as an alternative target quantity for diversity evaluation. We showed that existing Nyström and FKEA methods for approximating Vendi scores concentrate around this truncated population Vendi. An interesting future direction is to explore the relationship between other kernel approximation techniques and the truncated population Vendi. Also, a comprehensive analysis of the computational-statistical trade-offs involved in estimating the Vendi score is another relevant future direction.
Acknowledgements.
This work is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14209920, and is partially supported by CUHK Direct Research Grants with CUHK Project No. 4055164 and 4937054. Finally, the authors thank the anonymous reviewers for their thoughtful feedback and constructive suggestions.
Supplementary Material
@@ A.1 Proof of Theorem 1
To prove the theorem, we will use the following lemma followed from [Gross, 2011, Kohler and Lucchi, 2017].
Lemma 1 (Vector Bernstein Inequality [Gross, 2011, Kohler and Lucchi, 2017]).
@@
On the other hand, note that $\phi(x)\otimes\phi(x)$ is the vectorized version of rank-1 $\phi(x)\phi(x)^{\top}$, which shows that the above inequality is equivalent to the following where $\|\cdot\|_{\mathrm{HS}}$ denotes the Hilbert-Schmidt norm, which will simplify to the Frobenius norm in the finite dimension case,
$\mathbb{P}\Bigl{(}\Bigl{\|}\frac{1}{n}\sum_{i=1}^{n}\bigl{[}\phi(x_{i})\phi(x_{i})^{\top}\bigr{]}-\mathbb{E}_{x\sim P}\bigl{[}\phi(x)\phi(x)^{\top}\bigr{]}\Bigr{\|}_{\mathrm{HS}}\geq\epsilon\Bigr{)}\,\leq\,\exp\Bigl{(}\frac{8-n\epsilon^{2}}{32}\Bigr{)}$
@@
Subsequently, we can apply the Hoffman-Wielandt inequality which shows that for the sorted eigenvalue vectors of $C_{X}$ (denoted by $\widehat{\boldsymbol{\lambda}}_{n}$ in the theorem) and $\widetilde{C}_{X}$ (denoted by $\widetilde{\boldsymbol{\lambda}}$ in the theorem) we will have $\|\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\|_{2}\leq\|C_{X}-\widetilde{C}_{X}\|_{\mathrm{HS}}$, which together with the previous inequality leads to
@@ A.2 Proof of Corollary 2
The case of $\alpha=1$. We show that Theorem 1 on the concentration of the eigenvalues $\boldsymbol{\lambda}=[\lambda_{1},\ldots,\lambda_{d}]$ will further imply a concentration bound for the logarithm of Vendi-1 score. In the case of $\mathrm{Vendi}_{1}$ (when $\alpha\rightarrow 1^{+}$), the concentration bound will be formed for the logarithm of the Vendi score, i.e. the Von-Neumann entropy (denoted as $H_{\alpha}$):
$H_{1}(C_{X}):=H_{1}(\boldsymbol{\lambda})=\sum_{i=1}^{d}\widetilde{\lambda}_{i}\log\frac{1}{\widetilde{\lambda}_{i}}$
Theorem 1 shows that $\|\widehat{\boldsymbol{\lambda}}_{n}-\widetilde{\boldsymbol{\lambda}}\|_{2}\leq\sqrt{\frac{32\log(2/\delta)}{n}}$ with probability $1-\delta$. To convert this concentration bound to a bound on the order-1 entropy (for Vendi-1 score) difference $H_{1}({\widehat{C}_{n}})-H_{1}(C_{X})$, we leverage the following two lemmas:
Lemma 2.
For every $0\leq\alpha,\beta\leq 1$ such that $|\beta-\alpha|\leq\frac{1}{e}$, we have
$\Bigl{|}\alpha\log\frac{1}{\alpha}-\beta\log\frac{1}{\beta}\Bigr{|}\,\leq\,|\beta-\alpha|\log\frac{1}{|\beta-\alpha|}$
Proof. Let $c=|\alpha-\beta|$, where $c\in[0,\frac{1}{e}]$. Defining $g(z)=z\log(\frac{1}{z})$, the first-order optimality condition $g^{\prime}(z)=-\log(z)-1=0$ yields $\frac{1}{e}$ as the local maximum of $g(z)$. Therefore, there are three cases of placement of $\alpha$ and $\beta$ on the interval $[0,1]$: $\alpha$ and $\beta$ appear before maximum point, after maximum point or maximum point is between $\alpha$ and $\beta$. We show that regardless of the placement of $\alpha$ and $\beta$, the above inequality remains true.
  • Case 1: $\alpha,\beta\in[0,\frac{1}{e}]$. Note that $g^{\prime\prime}(z)=-\frac{1}{z}$. Since the second-order derivative is negative and the function $g$ is monotonically increasing within the interval $[0,\frac{1}{e}]$, the gap between $g(\alpha)$ and $g(\beta)$ is maximized when $\alpha^{*}=0$ and $\beta^{*}=c-\alpha^{*}=c$. This directly leads to the desired bound as follows:
    $\Bigl{|}\alpha\log\frac{1}{\alpha}-\beta\log\frac{1}{\beta}\Bigr{|}\,\leq\,\Bigl{|}\alpha^{*}\log\frac{1}{\alpha^{*}}-\beta^{*}\log\frac{1}{\beta^{*}}\Bigr{|}\,=\,\bigl{|}0\log 0-c\log\frac{1}{c}\bigr{|}\,\leq\,c\log\frac{1}{c}$
    Here, we use the standard limit $00\log 0=$ 0.
  • Case 2: $\alpha,\beta\in[\frac{1}{e},1]$. In this case, we note that $g$ is concave yet decreasing over $[\frac{1}{e},1]$, and so the gap between $g(\alpha)$ and $g(\beta)$ will be maximized when $\alpha^{*}=1-c$ and $\beta^{*}=1$. This leads to:
    $\Bigl{|}\alpha\log\frac{1}{\alpha}-\beta\log\frac{1}{\beta}\Bigr{|}\,\leq\,\Bigl{|}\alpha^{*}\log\frac{1}{\alpha^{*}}-\beta^{*}\log\frac{1}{\beta^{*}}\Bigr{|}\,=\,(1-c)\log\frac{1}{(1-c)}\leq c\log\frac{1}{c}$
    where the last inequality holds because $c\in[0,\frac{1}{e}]$, and if we define the function $h(c)=c\log\frac{1}{c}-(1-c)\log\frac{1}{1-c}$, then we have $h^{\prime}(c)=\log\frac{1}{c(1-c)}-2$, which is positive over $c\in[0,c_{0}]$ ($e^{-2}<c_{0}<e^{-1}$ is where $c_{0}(1-c_{0})=e^{-2}$), and then negative over $[c_{0},\frac{1}{e}]$, and hence $h(c)\geq\min\{h(0),h(1/e)\}=0$ for every $c\in[0,1/e]$.
  • Case 3: $\alpha\in[0,\frac{1}{e})$ and $\beta\in(\frac{1}{e},1]$. When $\alpha$ and $\beta$ lie on the opposite ends from the maximum point, the inequality becomes:
    $\bigl{|}\alpha\log\frac{1}{\alpha}-\beta\log\frac{1}{\beta}\bigr{|}\leq\mathrm{max}\Bigr{\{}\bigl{|}(1/e)\log\frac{1}{1/e}-\beta\log\frac{1}{\beta}\bigr{|},\bigr{|}\alpha\log\frac{1}{\alpha}-(1/e)\log\frac{1}{1/e}\bigl{|}\Bigl{\}}\leq c\log\frac{1}{c}$
    since we pick the side with the largest difference, this difference is upper bounded by either Case 1 or Case 2 because $\mathrm{max}\{|\frac{1}{e}-\beta|,|\alpha-\frac{1}{e}|\}<c$. Therefore, this case is upper-bounded by $c\log\frac{1}{c}$.
All the three cases of placement of $\alpha$ and $\beta$ are upper-bounded by $c\log\frac{1}{c}$; Therefore, the claim holds.
Lemma 3.
If $\|\mathbf{u}\|_{2}\leq\epsilon$ for $d$-dimensional vector $\mathbf{u}\geq\mathbf{0}$ where $\epsilon\leq\frac{1}{e}$, then we have
$\sum_{i=1}^{d}u_{i}\log\frac{1}{u_{i}}\leq\epsilon\sqrt{d}\log\frac{\sqrt{d}}{\epsilon}$
Proof. We prove the above inequality using the KKT conditions for the following maximization problem, representing a convex optimization problem,
$\max_{\mathbf{u}\in\mathbb{R}^{d}}\qquad\;\;\>\sum_{i=1}^{d}u_{i}\log(\frac{1}{u_{i}})$
$\text{\rm subject to}\quad u_{i}\geq 0,\;\;\text{\rm for all}\>i$
$\qquad\qquad\quad\sum_{i=1}^{d}u_{i}^{2}\leq\epsilon^{2}\;\;\ (\text{equivalent to }\|\mathbf{u}\|_{2}\leq\epsilon)$
In a concave maximization problem subject to convex constraints, any point that satisfies the KKT conditions is guaranteed to be a global optimum. Let us pick the following solution $\mathbf{u}^{*}=\frac{\epsilon}{\sqrt{d}}\mathbf{1}$ and slack variables $\lambda^{*}=\frac{\sqrt{d}}{2\epsilon}\bigr{(}\log(\frac{\sqrt{d}}{\epsilon})-1\bigl{)}$, $\forall_{i}\;\mu_{i}^{*}=0$. The Lagrangian of the above problem:
$L(\mathbf{u},\lambda,\mu_{1},\dots,\mu_{d})=\sum_{i=1}^{d}u_{i}\log(\frac{1}{u_{i}})+\lambda(\epsilon^{2}-\sum_{i=1}^{d}u_{i}^{2})-\sum_{i=1}^{d}\mu_{i}u_{i}$
  • Primal Feasibility. The solution $\mathbf{u^{*}}$ satisfies the primal feasibility, since $\epsilon^{2}-\sum_{i=1}^{d}(\frac{\epsilon}{\sqrt{d}})^{2}=\epsilon^{2}-d\frac{\epsilon^{2}}{d}=0$ and $\frac{\epsilon}{\sqrt{d}}\geq 0$.
  • Dual Feasibility. $\lambda^{*}\geq 0$ is feasible because of the assumption $\epsilon\leq\frac{1}{e}$ implying that $\frac{\sqrt{d}}{\epsilon}\geq e$ for every integer dimension $d\geq 1$. Note that this implies $\lambda^{*}=\frac{\sqrt{d}}{2\epsilon}\bigr{(}\log(\frac{\sqrt{d}}{\epsilon})-1\bigl{)}\geq 0$.
  • Complementary Slackness. Since $\lambda^{*}\bigr{(}\epsilon^{2}-\sum_{i=1}^{d}(\frac{\epsilon}{\sqrt{d}})^{2}\bigl{)}=\lambda^{*}\cdot 0=0$, the condition is satisfied.
  • Stationarity. The condition is satisfied as follows:
    $\frac{\partial}{\partial u_{i}}L(\mathbf{u^{*}})$ $=-\log(u_{i}^{*})-1-2\lambda^{*}u_{i}^{*}+\mu_{i}^{*}=-\log(\frac{\epsilon}{\sqrt{d}})-1-2\cdot\frac{\sqrt{d}}{2\epsilon}\bigr{(}-\log(\frac{\epsilon}{\sqrt{d}})-1\bigl{)}\cdot\frac{\epsilon}{\sqrt{d}}=0$
Since all KKT conditions are satisfied and sufficient for global optimality, $\mathbf{u}^{*}=\frac{\epsilon}{\sqrt{d}}\mathbf{1}$ is a global optimum of the specified concave maximization problem. We note that this result is also implied by the Schur-concavity property of entropy. Following this result, the specified objective is upper-bounded as follows:
$\sum_{i=1}^{d}u_{i}\log\frac{1}{u_{i}}\leq\epsilon\sqrt{d}\log\frac{\sqrt{d}}{\epsilon}$
Therefore, the lemma’s proof is complete.
Following the above lemmas, knowing that $\|\widehat{\lambda}_{n}-\widetilde{\lambda}\|_{2}\leq\sqrt{\frac{32\log(2/\delta)}{n}}$ from Theorem 1 and using the assumption $n\geq 32e^{2}\log(2/\delta)\approx 236.5\log(2/\delta)$ that ensures the upper-bound satisfies $\sqrt{\frac{32\log(2/\delta)}{n}}\leq\frac{1}{e}$, we can apply the above two lemmas to show that with probability $1-\delta$:
$\Bigl{|}H_{1}({\widehat{C}_{n}})-H_{1}(C_{X})\Bigr{|}=\Bigl{|}H_{1}(\widehat{\lambda}_{n})-H_{1}(\widetilde{\lambda})\Bigr{|}\leq\sqrt{\frac{8d\log(2/\delta)}{n}}\log\Bigl{(}\frac{nd}{32\log(2/\delta)}\Bigr{)}$
Note that under a kernel function with finite dimension $d$, the above bound will be $\mathcal{O}\Bigl{(}\sqrt{\frac{d}{n}}\log\bigl{(}nd\bigr{)}\Bigr{)}$.
The case of $21<\alpha<$ 2. Note that the inequality $\|v\|_{\alpha}\leq d^{\frac{2-\alpha}{2}}\|v\|_{2}$ holds for every $d$-dimensional vector $v\in\mathbb{R}^{d}$. Therefore, we can repeat the proof of Corollary 1 to show the following for every $21<\alpha<$ 2
A.3 Proof of Corollary 1
@@ A.4 Proof of Theorem 2
Proof. To prove the lemma, first note that $\mathbf{v}^{(t)}\in\Delta_{t}$, i.e. its first $t$ entries are non-negative and add up to $1$, and also its last $d-t$ entries are zero. Then, consider the projection problem discussed in the lemma:
@@
Then, since we know from the assumptions that $v_{i}\geq 0$ and $\sum_{i=1}^{t}v_{i}\leq 1$, the discussed $\mathbf{u}^{*}\in\mathbb{R}^{t}$ where $u^{*}_{i}=v_{i}+(1-S_{t})/t$ together with Lagrangian coefficients $\mu_{i}=0$ (for inequality constraint $u_{i}\geq 0$) and $\lambda=(1-S_{t})/t$ (for equality constraint) satisfy the KKT conditions. The primal and dual feasibility conditions as well as the complementary slackness clearly hold for these selection of primal and dual variables. Also, the KKT stationarity condition is satisfied as for every $i$ we have $u^{*}_{i}-v_{i}-\lambda-\mu_{i}=0$. Since the optimization problem is a convex optimization task with affine constraints, the KKT conditions are sufficient for optimaility which proves the lemma.
Based on the above lemma, the eigenvalues $\widehat{\boldsymbol{\lambda}}_{n}^{(t)}$ used to calculate the $t$-truncated Vendi score $\mathrm{Vendi}^{(t)}_{\alpha}(x_{1},\ldots,x_{n})$ are the projections of the top-$t$ eigenvalues in $\widehat{\boldsymbol{\lambda}}_{n}$ for the original score $\mathrm{Vendi}_{\alpha}(x_{1},\ldots,x_{n})$ onto the $t$-simplex subset of $\mathbb{R}^{d}$ according to the $\ell_{2}$-norm. Similarly, the eigenvalues $\widetilde{\boldsymbol{\lambda}}_{n}^{(t)}$ used to calculate the $t$-truncated population Vendi $\mathrm{Vendi}^{(t)}_{\alpha}(P_{X})$ are the projections of the top-$t$ eigenvalues in $\widetilde{\boldsymbol{\lambda}}$ for the original population Vendi $\mathrm{Vendi}_{\alpha}(P_{x})$ onto the $t$-simplex subset of $\mathbb{R}^{d}$.